One Integer is twice another. Their sum is 96. Find the integers.
Answer by mananth(16949)

let one integer =x
the other will be 2x
x+2x=96
3x=96
divide both sides by 3
3x/3 = 96/3
x=32 one integer
the other is 2x =2* 32=64
32,64